# NPLQCD collaboration

### v2

@@ Abstract
Understanding the behavior of dense hadronic matter is a central goal in nuclear physics as it governs the nature and dynamics of astrophysical objects such as supernovae and neutron stars. Because of the non-perturbative nature of quantum chromodynamics (QCD), little is known rigorously about hadronic matter in these extreme conditions. Here, lattice QCD calculations are used to compute thermodynamic quantities and the equation of state of QCD over a wide range of isospin chemical potentials. Agreement is seen with chiral perturbation theory predictions when the chemical potential is small. Comparison to perturbative QCD calculations at large chemical potential allows for an estimate of the gap in the superconducting phase, and this quantity is seen to agree with perturbative determinations. Since the partition function for an isospin chemical potential, $\mu_{I}$, bounds the partition function for a baryon chemical potential $\mu_{B}=3\mu_{I}/2$, these calculations also provide rigorous non-perturbative QCD bounds on the symmetric nuclear matter equation of state over a wide range of baryon densities for the first time.
@@
In this work, the first non-perturbative QCD constraint on the nuclear EoS with complete quantification of systematic uncertainties is presented. These calculations build upon the proof-of-principle, single lattice spacing results of Ref. Abbott:2023coj with improved methodology, increased statistical precision, an extrapolation to the continuum limit, and an interpolation to the physical quark masses, enabling a systematically-controlled result to be achieved for the first time. The pressure and other thermodynamic properties of low-temperature isospin-dense matter are determined over a wide range of densities and chemical potentials, spanning all scales from hadronic to perturbatively-coupled. At small values of the isospin chemical potential, $\mu_{I}$, the results are found to agree with chiral perturbation theory ($\chi$PT) son_qcd_2001 ; Kogut:2001id at next-to-leading order (NLO) Adhikari:2019zaj ; Andersen:2023ofv . At large $\mu_{I}$, the results are seen to agree with pQCD with pairing contributions Fujimoto:2023mvc . The comparison of next-to-next-to-leading-order (NNLO) pQCD predictions Freedman:1976xs ; Freedman:1976dm ; Freedman:1976ub ; Baluni:1977ms ; Kurkela:2009gj ; Fujimoto:2023mvc for the pressure (partial next-to-next-to-next-to-leading-order results are also available Gorda:2021znl ) with the continuum limit of the LQCD calculations provides a determination of the superconducting gap as a function of $\mu_{I}$. This is seen to agree with the leading-order perturbative calculation of the pairing gap Fujimoto:2023mvc , but is more precise. The speed of sound in isospin-dense matter is also seen to significantly exceed the conformal limit of $c_{s}^{2}/c^{2}\leq 1/3$ over a wide range of $\mu_{I}$. A Bayesian model mixing approach that combines $\chi$PT, LQCD and pQCD provides a determination of the zero-temperature EoS for isospin-dense QCD matter valid at all values of the isospin chemical potential for the first time.
From simple path-integral relations Cohen:2003ut ; Cohen:2004qp ; Fujimoto:2023unl ; Moore:2023glb , the determination of the pressure in isospin-dense matter provides a non-perturbative, model-independent bound on the pressure of isospin-symmetric QCD matter at nonzero baryon chemical potential and hence on the nuclear EoS. The current results therefore provide a systematically-controlled QCD bound at all densities, and the impact on nuclear phenomenology is briefly discussed.
@@
to determine the isospin chemical potential from the extracted energies. Other studies have added the isospin chemical potential directly to the QCD action Kogut:2002tm ; Kogut:2002zg ; Brandt:2017oyy ; Bornyakov:2021mfj ; Brandt:2022hwy . The primary advantage of the method used here in comparison to the canonical approach is that $\mu_{I}$ enters as an input to the calculation of thermodynamic quantities rather than being derived from the isospin charge of the LQCD data and therefore is not subject to statistical and systematic uncertainties. A more detailed comparison to the approach of Ref. Abbott:2023coj is discussed in the Supplementary Material
@@
By the monotonicity of the logarithm, the above inequality directly translates into an inequality between the pressures of the two media as a function of the energy density. Consequently, the isospin-dense EoS bounds the EoS for symmetric nuclear matter. At large values of the quark chemical potentials, where pQCD is valid, this bound becomes tight as differences between the partition functions enter only at ${\cal O}(\alpha_{s}^{k})$ for $k\geq 3$ Moore:2023glb . This bound was explored in Ref. Fujimoto:2023unl based on the previous lattice QCD results Abbott:2023coj at a single lattice-spacing and unphysical quark masses. Here, Fig. 4 presents updated bounds based on the continuum limit lattice QCD results at the physical quark masses, $\chi$PT, and perturbative QCD through the GP-model. While the bounds from isospin-dense matter do not significantly constrain phenomenological nuclear equations of state within the uncertainties that are typically presented Annala:2023cwx ; Semposki:2024vnp ; Han:2022rug ; Jiang:2022tps ; Raaijmakers:2019dks ; Huth:2020ozf ; vanDalen:2005sk ; Typel:2005ba ; Kolomeitsev:2004ff ; Liu:2001iz ; Leonhardt:2019fua , the bounds are independent of modeling uncertainties that enter the nuclear EoS in the regions that are unconstrained by nuclear structure or pQCD calculations.
@@ Acknowledgements.
We are grateful to Yuki Fujimoto and Sanjay Reddy for discussions. The ensembles used in this work were generated through the combined efforts of the JLab, William and Mary, Los Alamos, and MIT groups. We particularly thank Balint Joó for assistance with the generation of the gauge configurations and quark propagators used in this work. The calculations were performed using an allocation from the Innovative and Novel Computational Impact on Theory and Experiment (INCITE) program using the resources of the Oak Ridge Leadership Computing Facility located in the Oak Ridge National Laboratory, which is supported by the Office of Science of the Department of Energy under Contract DE-AC05-00OR22725. This research also used resources of the National Energy Research Scientific Computing Center (NERSC), a U.S. Department of Energy Office of Science User Facility located at Lawrence Berkeley National Laboratory, operated under Contract No. DE-AC02-05CH11231. We acknowledge USQCD computing allocations and PRACE for awarding us access to Marconi100 at CINECA, Italy. This work is supported by the National Science Foundation under Cooperative Agreement PHY-2019786 (The NSF AI Institute for Artificial Intelligence and Fundamental Interactions, http://iaifi.org/) and by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics under grant Contract Number DE-SC0011090. RA, WD and PES are also supported by the U.S. Department of Energy SciDAC5 award DE-SC0023116. FRL acknowledges financial support from the Mauricio and Carlota Botton Fellowship. MI is partially supported by the Quantum Science Center (QSC), a National Quantum Information Science Research Center of the U.S. Department of Energy. PES is also supported by the U.S. DOE Early Career Award DE-SC0021006. AP and RP acknowledge support from Grant CEX2019-000918-M and the project PID2020-118758GB-I00, financed by the Spanish MCIN/ AEI/10.13039/501100011033/, and from the EU STRONG-2020 project under the program H2020-INFRAIA-2018-1 grant agreement no. 824093. This manuscript has been authored by Fermi Research Alliance, LLC under Contract No. DE-AC02-07CH11359 with the U.S. Department of Energy, Office of Science, Office of High Energy Physics. This work made use of Chroma Edwards:2004sx , QDPJIT Wint1405:Framework , QUDA Clark:2010 ; Clark:2016rdz , JAX jax2018github , NumPy numpy , SciPy 2020SciPy-NMeth , matplotlib Hunter:2007 and HDF5 The_HDF_Group_Hierarchical_Data_Format .

### v3

@@ Abstract
Understanding the behavior of dense hadronic matter is a central goal in nuclear physics as it governs the nature and dynamics of astrophysical objects such as supernovae and neutron stars. Because of the non-perturbative nature of quantum chromodynamics (QCD), little is known rigorously about hadronic matter in these extreme conditions. Here, lattice QCD calculations are used to compute thermodynamic quantities and the equation of state of QCD over a wide range of isospin chemical potentials with controlled systematic uncertainties. Agreement is seen with chiral perturbation theory when the chemical potential is small. Comparison to perturbative QCD at large chemical potential allows for an estimate of the gap in the superconducting phase, and this quantity is seen to agree with perturbative determinations. Since the partition function for an isospin chemical potential, $\mu_{I}$, bounds the partition function for a baryon chemical potential $\mu_{B}=3\mu_{I}/2$, these calculations also provide rigorous non-perturbative QCD bounds on the symmetric nuclear matter equation of state over a wide range of baryon densities for the first time.
@@
In this work, the first non-perturbative QCD constraint on the nuclear EoS with complete quantification of systematic uncertainties is presented. These calculations build upon the proof-of-principle, single lattice spacing results of Ref. Abbott:2023coj with improved methodology, increased statistical precision, an extrapolation to the continuum limit, and an interpolation to the physical quark masses, enabling a systematically-controlled result to be achieved for the first time. The pressure and other thermodynamic properties of low-temperature isospin-dense matter are determined over a wide range of densities and chemical potentials, spanning all scales from hadronic to perturbatively-coupled. At small values of the isospin chemical potential, $\mu_{I}$, the results agree with chiral perturbation theory ($\chi$PT) son_qcd_2001 ; Kogut:2001id at next-to-leading order (NLO) Adhikari:2019zaj ; Andersen:2023ofv . At large $\mu_{I}$, the results are seen to agree with pQCD with pairing contributions Fujimoto:2023mvc . The comparison of next-to-next-to-leading-order (NNLO) pQCD predictions Freedman:1976xs ; Freedman:1976dm ; Freedman:1976ub ; Baluni:1977ms ; Kurkela:2009gj ; Fujimoto:2023mvc for the pressure (partial next-to-next-to-next-to-leading-order results are also available Gorda:2021znl ) with the continuum limit of the LQCD calculations provides a determination of the superconducting gap as a function of $\mu_{I}$. This is seen to agree with the leading-order perturbative calculation of the pairing gap Fujimoto:2023mvc , but is more precise. The speed of sound in isospin-dense matter is also seen to significantly exceed the conformal limit of $c_{s}^{2}/c^{2}\leq 1/3$ over a wide range of $\mu_{I}$. A Bayesian model mixing approach that combines $\chi$PT, LQCD and pQCD provides a determination of the zero-temperature EoS for isospin-dense QCD matter valid at all values of the isospin chemical potential.
These results provide constraints for phenomenological frameworks seeking to describe the QCD phase diagram MUSES:2023hyz , and from simple path-integral relations Cohen:2003ut ; Cohen:2004qp ; Fujimoto:2023unl ; Moore:2023glb , the determination of the pressure in isospin-dense matter provides a non-perturbative, model-independent bound on the pressure of isospin-symmetric QCD matter at nonzero baryon chemical potential and hence on the nuclear EoS. The current results therefore provide a systematically-controlled QCD bound at all densities, and the impact on nuclear phenomenology is briefly discussed.
@@
to determine the isospin chemical potential from the extracted energies. Other studies have added the isospin chemical potential directly to the QCD action Kogut:2002tm ; Kogut:2002zg ; Brandt:2017oyy ; Bornyakov:2021mfj ; Brandt:2022hwy . These studies probe isospin dense QCD at $\mu_{I}\lesssim 2$ and focus primarily on nonzero temperature. The results in this work are consistent with the low-temperature results in those works, but span a larger range of chemical potentials.
The primary advantage of the method used here in comparison to the method of Ref. Abbott:2023coj is that $\mu_{I}$ enters as an input to the calculation of thermodynamic quantities rather than being derived from the isospin charge of the LQCD data and therefore is not subject to statistical and systematic uncertainties. A more detailed comparison to the approach of Ref. Abbott:2023coj is discussed in the Supplementary Material
@@
By the monotonicity of the logarithm, the above inequality directly translates into an inequality between the pressures of the two media as a function of the energy density. Consequently, the isospin-dense EoS bounds the EoS for symmetric nuclear matter. At large values of the quark chemical potentials, where pQCD is valid, this bound becomes tight as differences between the partition functions enter only at ${\cal O}(\alpha_{s}^{k})$ for $k\geq 3$ Moore:2023glb . This bound was explored in Ref. Fujimoto:2023unl based on the previous lattice QCD results Abbott:2023coj at a single lattice-spacing and unphysical quark masses. Here, Fig. 4 presents updated bounds based on the continuum limit lattice QCD results at the physical quark masses, $\chi$PT, and perturbative QCD through the GP-model. While the bounds from isospin-dense matter do not significantly constrain phenomenological nuclear equations of state within the uncertainties that are typically presented Annala:2023cwx ; Semposki:2024vnp ; Han:2022rug ; Jiang:2022tps ; Raaijmakers:2019dks ; Huth:2020ozf ; vanDalen:2005sk ; Typel:2005ba ; Kolomeitsev:2004ff ; Liu:2001iz ; Leonhardt:2019fua , the bounds are independent of modeling uncertainties that enter the nuclear EoS in the regions that are unconstrained by nuclear structure or pQCD calculations. GP models without lattice QCD constraints result in significantly larger uncertainties in the position of the bound, in particular in the lower right corner of the red band of Fig. 4.
@@ Acknowledgements.
We are grateful to Yuki Fujimoto and Sanjay Reddy for discussions. The ensembles used in this work were generated through the combined efforts of the JLab, William and Mary, Los Alamos, and MIT groups. We particularly thank Balint Joó for assistance with the generation of the gauge configurations and quark propagators used in this work. The calculations were performed using an allocation from the Innovative and Novel Computational Impact on Theory and Experiment (INCITE) program using the resources of the Oak Ridge Leadership Computing Facility located in the Oak Ridge National Laboratory, which is supported by the Office of Science of the Department of Energy under Contract DE-AC05-00OR22725. This research also used resources of the National Energy Research Scientific Computing Center (NERSC), a U.S. Department of Energy Office of Science User Facility located at Lawrence Berkeley National Laboratory, operated under Contract No. DE-AC02-05CH11231. We acknowledge USQCD computing allocations and PRACE for awarding us access to Marconi100 at CINECA, Italy. This work is supported by the National Science Foundation under Cooperative Agreement PHY-2019786 (The NSF AI Institute for Artificial Intelligence and Fundamental Interactions, http://iaifi.org/) and by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics under grant Contract Number DE-SC0011090. RA, WD and PES are also supported by the U.S. Department of Energy SciDAC5 award DE-SC0023116. RA was also partially supported by the High Energy Physics Computing Traineeship for Lattice Gauge Theory (DE-SC0024053). FRL acknowledges financial support from the Mauricio and Carlota Botton Fellowship. MI is partially supported by the Quantum Science Center (QSC), a National Quantum Information Science Research Center of the U.S. Department of Energy. PES is also supported by the U.S. DOE Early Career Award DE-SC0021006. AP and RP acknowledge support from Grant CEX2019-000918-M and the project PID2020-118758GB-I00, financed by the Spanish MCIN/ AEI/10.13039/501100011033/, and from the EU STRONG-2020 project under the program H2020-INFRAIA-2018-1 grant agreement no. 824093. This manuscript has been authored by Fermi Research Alliance, LLC under Contract No. DE-AC02-07CH11359 with the U.S. Department of Energy, Office of Science, Office of High Energy Physics. This work made use of Chroma Edwards:2004sx , QDPJIT Wint1405:Framework , QUDA Clark:2010 ; Clark:2016rdz , JAX jax2018github , NumPy numpy , SciPy 2020SciPy-NMeth , matplotlib Hunter:2007 and HDF5 The_HDF_Group_Hierarchical_Data_Format .